# references.bib

### v4

@@ 3.1 The groupoid approach for H-elliptic operators
Let $M$ be a smooth closed filtered manifold. We consider the global groupoid structure which constructed over $[0,1]^{2}$ with units in $M\times[0,1]^{2}$ as follows:
@@
By the same way, the groupoids constructed above can be attached by the diagram
@@
At $s=1$, the attached groupoid is the Heisenberg tangent groupoid, or is called the holonomy groupoid of the adiabatic foliation $\mathfrak{a}H$ associated with $H$. Further information on holonomy groupoid and adiabatic foliation is available in [androulidakis2022pseudodifferential]
We apply the restriction of the global groupoid $C^{\ast}$-algebra to each side. Then, as discussed in Section 1.7, for a family of $C^{\ast}$-algebras over $[0,1]$, with $B$ fibered over $(0,1]$ and $A$ over 00, we obtain an exact sequence that induces a map in $K$-theory or $KK$-theory that
@@
Note that the $\mu_{s=0}$ is induced by the attaching between $gr(H)^{\ast}$ and $T^{\ast}M$ facilitated by $\delta_{t}^{-1}$. Subsequently, we explore the map between the $KK$-groups of $C^{\ast}$-algebras induced by the isomorphism from the bundle $TM$ to $gr(H)$ via orthogonal decomposition—such an isomorphism is called a splitting isomorphism ([van2017tangent]). The induced map in $K$-theory remains invariant regardless of the selected isomorphism, leading to a concurrence of these two induced maps through direct computation.
@@
The index map $\text{Ind}_{H^{\bullet}}:K_{0}(Gr(H))\to\mathbb{Z}$ can be evaluated as
@@
where $\mu_{t=0}^{-1}:K_{0}(C^{\ast}(Gr(H)))\to K^{0}(gr(H)^{\ast})$ representing Connes-Thom’s isomorphism. This constitutes the index theorem for H-elliptic differential operators.
An alternative formulation of this index theorem will be given in the next section and be proved in Chapter 5 using the $KK$-theoretic argument.
This section derives the index theorem for H-elliptic differential operators from the $KK$-theoretic perspective, expressed as a $KK$-product representation. We focus on the index of operators in $KK_{0}$, the $KK_{1}$ index for self-adjoint operators on non-graded bundles can be analogously analyzed. The H-elliptic operators can be analyzed using Kasparov’s $KK$-theory.
In the $G$-invariant cases, the Lie group $G$ acts on $TM$ induced by the $G$-action on $M$. We assume that $G$ preserves the H-filtration structure. Then, we have
@@ 3.2 Index theorem for H-elliptic operators from $KK$-theoretic viewpoint
For a general H-elliptic differential operator $P_{0}$, we consider the bounded transform
$P=\begin{pmatrix}0&P_{0}^{\ast}/(1+P_{0}P_{0}^{\ast})\\ P_{0}/(1+P_{0}^{\ast}P_{0})&0\end{pmatrix},$
which is a $G$-invariant Fredholm operator of order 00. $[a,P]$ and $a(1-P^{2})$ are operators with negative order symbols, so they are both compact. Hence, we can define that
@@
The index class $[P_{0}]$ of $P_{0}$ is defined to be the $KK$-class $[(L^{2}(E_{0}\oplus E_{1}),P)]\in KK_{G}(\mathbb{C},\mathbb{C})\cong R(G)$.
@@
Apply $\mathcal{G}=Gr(H)$ in Proposition 1.7.5, the Connes-Thom’s isomorphism can be represented by a $KK$-product with a certain element $[\mathcal{E}_{0}]\in KK(C^{\ast}(Gr(H)),$ $C_{0}(gr(H)^{\ast}))$. Let
@@
The analytical index map $\text{Ind}_{H^{\bullet}}$ for H-elliptic operators can be expressed by $KK$-products as
@@
The Kasparov product with $[\mathcal{E}]$ facilitates an isomorphism between $KK(\mathbb{C},C^{\ast}(Gr(H))$ and $KK(\mathbb{C},C_{0}(T^{\ast}M))$, which represents the composition of the Connes-Thom’s isomorphism and the induced isomorphism of $K^{0}$ between $T^{\ast}M$ and $gr(H)^{\ast}$, i.e.,
@@
The Kasparov product with $[\mathcal{D}_{M}]$ gives the ordinary analytical index map, i.e.,
@@
$[\mathcal{E}]$ represent a $G$-invariant $KK$-class in $KK_{G}(C^{\ast}(Gr(H)),C_{0}(T^{\ast}M))$. The Kasparov product with $[\mathcal{E}]$ provides an isomorphism between $KK_{G}(\mathbb{C},$ $C^{\ast}(Gr(H))$ and $KK_{G}(\mathbb{C},C_{0}(T^{\ast}M))$.
Thus, the index theorem for H-elliptic operators in $KK$-theory can be expressed as
@@
Let $M$ be a compact filtered Riemannian manifold and $G$ be a Lie group, which isometrically acts on $M$ and preserves the filtered structure. If $P_{0}$ be an H-elliptic $G$-invariant differential operator on $M$, let
$P=\begin{pmatrix}0&P_{0}^{\ast}/(1+P_{0}P_{0}^{\ast})\\ P_{0}/(1+P_{0}^{\ast}P_{0})&0\end{pmatrix}$
be the bounded transform of $P_{0}$ of order 00, then we have
@@
$KK_{G}(\mathbb{C},C^{\ast}(Gr(H))$ $\otimes_{C^{\ast}(Gr(H))}KK_{G}(C^{\ast}(Gr(H)),C_{0}(T^{\ast}M))$
$\otimes_{C_{0}(T^{\ast}M)}KK_{G}(C_{0}(T^{\ast}M),\mathbb{C}).$
The proof will be given in chapter 5.
Later, we will use the following lemma in the proof of Theorem 3.2.6 and the main Theorem 5.2.2.
Lemma 3.2.7.
Let $P$ be the bounded transform of $P_{0}$ as in Theorem 3.2.6. For the multiplier $\sigma_{H}(P)$ of $C^{\ast}(Gr(H))$, representing an element of $KK_{G}(\mathbb{C},C^{\ast}(Gr(H)))$, the product with $[\mathcal{E}]$ induces the transformation as the following three steps:
  • (i)
    Freezing the coefficient of the multiplier $\sigma_{H}(P)(x)=P_{x}$ to an element in $\mathcal{U}(gr(H)_{x})$ for each $x\in M$, which represents a differential operator on $C^{\ast}(gr(H)_{x})$ as $gr(H)_{x}$ is a vector space.
  • (ii)
    Choose a splitting isomorphism between the bundle $gr(H)$ and $TM$ (herein, we presume it to be the orthonormal decomposition in Definition 4.1.1), the differential operator family on $C^{\ast}(gr(H)_{x})$ can be transformed into a family of operators $P^{\prime}=\{P^{\prime}_{x}\}_{x\in M}$ acting on $C^{\ast}(T_{x}M)$ for each $x\in M$.
  • (iii)
    Take the Fourier transform from $C^{\ast}(TM)$ to $C_{0}(T^{\ast}M)$, the family of operators $P^{\prime}$ will be transformed into the multiplication operator $\sigma^{\prime}(P)$ on $C_{0}(T^{\ast}M)$ (If $P$ act on bundle $E$, we need to tensor product endomorphisms on $E$), which represent an element in $K^{0}(T^{\ast}M)$
This procedure yields a multiplicative operator $\sigma^{\prime}(P)(x,\xi)$ on $C_{0}(TM)$, which is essentially the Fourier transform of the freezing coefficient operators of the H-principal component of the original operator $P$.
Proof.
First, let $P$ be the H-elliptic $G$-invariant differential operator $P_{0}$.
The product with $[\mathcal{E}]$ is the composition of isomorphisms $\mu_{t=0}^{-1}$ and $\mu_{s=0}$. The first one is induced from the adiabatic groupoid for $Gr(H)_{x}$ as in Proposition 1.7.5. Let $P_{0,x}$ be the freezing coefficient of $P_{x}$ to $e\in Gr(H)_{x}$, by the arguments in [van2015groupoid], we see that on this adiabatic groupoid, the family $\mathbb{P}_{t}$ defined by
$\mathbb{P}_{t}=\begin{cases}t^{m}P_{x},&\quad t\neq 0\\ P_{0,x},&\quad t=0\\ \end{cases}$
is smooth. Thus the Connes-Thom isomorphism $\mu_{t=0}$ transforms the $KK$-class $[P_{0,x}]$ to $[P_{x}]$. So the inverse map $\mu_{t=0}^{-1}$ transforms from $[P_{x}]$ to $[P_{0,x}]$. The isomorphism $\mu_{s=0}$ is induced by the splitting isomorphism from $T_{x}M$ to $gr(H)_{x}$ so it transforms to an operator $P^{\prime}_{x}$ on $T_{x}M$. As a result, we attach these operators over $x\in M$ and obtain a family of operators $P^{\prime}=P^{\prime}_{x}$ such that the product with $[\mathcal{E}]$ transforms $[\sigma_{H}(P)]\in KK_{G}(\mathbb{C},C^{\ast}(Gr(H)))$ to $[P^{\prime}]\in KK_{G}(\mathbb{C},C^{\ast}(TM))$. Take the Fourier transform from $C^{\ast}(TM)$ to $C_{0}(T^{\ast}M)$ we obtain $\sigma^{\prime}(P)(x,\xi)$.
Note that for a differential operator $P=\sum_{\lvert\alpha\rvert_{H}\leq m}c_{\alpha}X^{\alpha}$ on $M$, the term $P_{0,x}$ in above process is $P_{0,x}=\sum_{\lvert\alpha\rvert_{H}=m}c_{\alpha}[X]_{x}^{\alpha}\in\mathcal{U}(gr(H)_{x})$, and the $KK$-class only depend on the principal part, so we can change the splitting isomorphism to the map $[X]_{x}\in gr(H)_{x}\mapsto X_{x}\in T_{x}M$. Thus we can easily see that $P^{\prime}_{x}$ can be taken to be the freezing coefficient of the H-principal component of the original operator $P$.
Take the functional calculus by
$T(a)=\begin{pmatrix}0&a^{\ast}/(1+aa^{\ast})\\ a/(1+a^{\ast}a)&0\end{pmatrix}$
on both $P$ and $\sigma^{\prime}(P)$, we see that the same result holds replacing $P$ by its bounded transformation. ∎

### v5

@@ 3.1 The groupoid approach for H-elliptic operators
In [mohsen2022index], Mohsen constructed a diagram of groupoids over $[0,1]^{2}$, such that the groupoid structure on each side independently forms a Lie groupoid structure. A $C^{\ast}$-algebra is defined on this groupoid, analyzed, and integrated to deduce the index theorem for H-elliptic differential operators.
Let $M$ be a smooth closed filtered manifold. We consider the global groupoid structure which is constructed over $[0,1]^{2}$ with units in $M\times[0,1]^{2}$ as follows:
@@
In the same way, the groupoids constructed above can be attached by the diagram
@@
At $s=1$, the attached groupoid is the Heisenberg tangent groupoid, or is called the holonomy groupoid of the adiabatic foliation $\mathfrak{a}H$ associated with $H$. Further information on holonomy groupoid and adiabatic foliation is available in [androulidakis2022pseudodifferential].
We apply the restriction of the global $C^{\ast}$-algebra of the constructed groupoid over $[0,1]^{2}$ to each side of the square. Then, as discussed in Section 1.8, for a family of $C^{\ast}$-algebras over $[0,1]$, with $B$ fibered over $(0,1]$ and $A$ over 00, we obtain an exact sequence that induces a map in $K$-theory or $KK$-theory that
@@
Note that the $\mu_{s=0}$ is induced by the attaching between $gr(H)^{\ast}$ and $T^{\ast}M$ facilitated by $\delta_{t}^{-1}$. Subsequently, we explore the map between the $KK$-groups of $C^{\ast}$-algebras is induced by the isomorphism from the bundle $TM$ to $gr(H)$ via orthogonal decomposition—such an isomorphism is called a splitting isomorphism ([van2017tangent]). The induced map in $K$-theory remains invariant regardless of the selected isomorphism, leading to the consistency of these two induced maps through direct computation.
@@
The index map $\text{Ind}_{H^{\bullet}}:K_{0}(C^{\ast}(Gr(H))\to\mathbb{Z}$ can be evaluated as
@@
where $\mu_{t=0}^{-1}:K_{0}(C^{\ast}(Gr(H)))\to K^{0}(gr(H)^{\ast})$ representing Connes-Thom’s isomorphism.
This constitutes the index theorem for H-elliptic differential operators. An alternative formulation of this index theorem will be given in the next section.
3.2 Mohsen’s Index theorem for H-elliptic operators from $KK$-theoretic viewpoint
H-elliptic operators can also be analyzed using Kasparov’s $KK$-theory. In this section, we transform the index theorem for H-elliptic differential operators (Theorem 3.1.3), to be expressed as a $KK$-product representation.
In the $G$-invariant cases, there is a Lie group $G$ action on $TM$ induced by the $G$-action on $M$. We assume that $G$ preserves the H-filtration structure. Then, we have
@@ 3.2 Mohsen’s Index theorem for H-elliptic operators from $KK$-theoretic viewpoint
For a general $G$-invariant H-elliptic differential operator $P_{0}$, we consider the bounded transform
$P=\begin{pmatrix}0&P_{0}^{\ast}/(1+P_{0}P_{0}^{\ast})^{1/2}\\ P_{0}/(1+P_{0}^{\ast}P_{0})^{1/2}&0\end{pmatrix},$
which is a $G$-invariant Fredholm operator of order 00, and for $a\in C_{0}(M)$, $[a,P]$ and $a(1-P^{2})$ are operators with negative order symbols, so they are both compact. Hence, we can define that
@@
The index class $[P_{0}]$ of $P_{0}$ is defined to be the $KK$-class $[(L^{2}(E_{0}\oplus E_{1}),P)]\in K^{0}_{G}(C_{0}(M))$.
@@
The commutative diagram in Section 3.1 can be obtained for $KK((C_{0}(M)),\cdot)$ using a similar argument, then we have the following diagram:
${KK(C_{0}(M),C_{0}(gr(H)^{\ast}))}$${KK(C_{0}(M),C_{0}(T^{\ast}M))}$${KK(C_{0}(M),C^{\ast}(Gr(H)))}$${KK(C_{0}(M),\mathbb{C}).}$$\scriptstyle{\mu_{t=0}}$$\scriptstyle{\mu_{s=0}}$$\scriptstyle{\mu_{t=1}}$$\scriptstyle{\mu_{s=1}}$
The Connes-Thom’s isomorphism can be represented by a $KK$-product with a certain element $[\mathcal{E}_{0}]\in KK(C^{\ast}(Gr(H)),C_{0}(gr(H)^{\ast}))$. Let
@@
The analytical index map $\text{Ind}_{H^{\bullet}}$ for H-elliptic operators in Proposition 1.7.4 can be expressed by $KK$-products as
@@
The Kasparov product with $[\mathcal{E}]$ facilitates an isomorphism between $KK(C_{0}(M),C^{\ast}(Gr(H))$ and $KK(C_{0}(M),C_{0}(T^{\ast}M))$, which represents the composition of the Connes-Thom’s isomorphism and the induced isomorphism of $K^{0}$ between $T^{\ast}M$ and $gr(H)^{\ast}$, i.e.,
@@
By Theorem 2.1.4, the Kasparov product with $[\mathcal{D}_{M}]$ gives the ordinary analytical index map, i.e.,
@@
$[\mathcal{E}]$ represent a $G$-invariant $KK$-class in $KK^{G}(C^{\ast}(Gr(H)),C_{0}(T^{\ast}M))$. The Kasparov product with $[\mathcal{E}]$ provides an isomorphism between $KK^{G}(C_{0}(M),C^{\ast}(Gr(H))$ and $KK^{G}(C_{0}(M),C_{0}(T^{\ast}M))$.
Thus, we can give an equivalent statement of the index theorem for H-elliptic operators (Theorem 3.1.3) in $KK$-theory as follows:
@@
Let $M$ be a complete filtered Riemannian manifold and $G$ be a locally compact Lie group, which isometrically acts on $M$ and preserves the filtered structure. Let $P$ is an H-elliptic $G$-invariant differential operator on $M$ of order 00, then we have
@@
$KK^{G}(C_{0}(M),C^{\ast}(Gr(H))$ $\otimes_{C^{\ast}(Gr(H))}KK^{G}(C^{\ast}(Gr(H)),C_{0}(T^{\ast}M))$
$\otimes_{C_{0}(T^{\ast}M)}KK^{G}(C_{0}(T^{\ast}M),\mathbb{C}).$
Remark 3.2.7.
For the multiplier $\sigma_{H}(P)$ of $C^{\ast}(Gr(H))$, it represents an element of $KK(\mathbb{C},C^{\ast}(Gr(H)))$, its $KK$-product with $[\mathcal{E}]$ lies in $K^{0}(T^{\ast}M)$ and is represented by an elliptic symbol $\sigma^{\prime}(P)(x,\xi)$ on $C_{0}(TM)$. The construction from $P$ to $\sigma^{\prime}(P)$ is natural, however, it is quite difficult to give an explicit formula for this elliptic symbol from the H-elliptic operator $P$. Even in the simplest case that $M$ is a contact manifold, the construction of this elliptic symbol is complicated, see [baum2014k]. That is why it is difficult to analyze the H-elliptic index theory explicitly in $KK$-theory.
In this section, we will see the index theory for hypoelliptic operators discussed in [kasparov2024coarsepseudodifferentialcalculusindex]. Note that in our case that $M$ is a filtered manifold, we always consider H-elliptic operators. First, we introduce some basic $KK$-elements.
Cover $M$ with a smooth family of small balls $\{U_{x}\}$ with centers $x\in M$ of radius $r_{x}$. Assume that the radii $r_{x}$ vary smoothly over $M$. This family of balls actually defines an open neighborhood $U$ of the diagonal in $M\times M$. We assume that both coordinate projections of the closure of $U$ into $M$ are proper maps.
Definition 3.3.1.
Recall that in Definition 2.2.4, we have defined the algebra $Cl_{\tau}(M)$. Define the radical covector field $\Theta_{x}(y)=\rho(x,y)d_{y}\rho(x,y)/r_{x}$ with $y\in U_{x}$. Consider $\Theta_{x}(y)$ as an element of the Clifford algebra fiber of $Cl_{\tau}(U_{x})$ at the point $y\in U_{x}$. By definition $\Theta_{x}^{2}-1\in C_{0}(U_{x})\subset Cl_{\tau}(U_{x})$, so globally over $M$, the family of Clifford multiplications by covector fields $\Theta_{x}$ defines an element $[\Theta_{M}]\in KK^{G}(C_{0}(M),C_{0}(U)\cdot C_{0}(M)\otimes Cl_{\tau}(M))$, and consequently, an element of $KK^{G}(C_{0}(M),C_{0}(M)\otimes Cl_{\tau}(M))$. It is called the local dual Dirac element.
Definition 3.3.2.
The Dirac element $[d_{M}]\in K^{0}_{G}(Cl_{\tau}(M))$ is defined as follows. Let $H=L^{2}(\Lambda^{\ast}(M))$ be the Hilbert space of complex-valued $L^{2}$-forms on $M$ graded by the even-odd form decomposition. The homomorphism $Cl_{\tau}(M)\to B(H)$ is given on (real) covector fields by the Clifford multiplication operators $v\mapsto\text{ext}(v)+\text{int}(v)$. The (unbounded) operator $d_{M}$ is the operator of the exterior derivation on $H$. The operator $D_{M}=d_{M}+d^{\ast}_{M}$ is essentially self-adjoint. We define $[d_{M}]$ as the pair $(H,D_{M}(1+D_{M}^{2})^{-1/2})$.
Theorem 3.3.3 ([kasparov1988equivariant], Theorem 4.8).
Let $M$ be a $G$-manifold. We have
$[\Theta_{M}]\otimes_{Cl_{\tau}(M)}[d_{M}]=1_{M}\in KK^{G}(C_{0}(M),C_{0}(M)).$
From 5.1 in [kasparov1988equivariant], for a Lie group $\mathcal{G}$ with maximal compact subgroup $\mathcal{K}$, there exists an element $\eta_{\mathcal{G}/\mathcal{K}}\in K^{\mathcal{G}}_{0}(Cl_{\tau}(\mathcal{G}/\mathcal{K}))$ which satisfies the following two equivalent conditions:
$[d_{\mathcal{G}/\mathcal{K}}]\otimes_{\mathbb{C}}[\eta_{\mathcal{G}/\mathcal{K}}]=1_{Cl_{\tau}(\mathcal{G}/\mathcal{K})}\in KK^{\mathcal{G}}(Cl_{\tau}(\mathcal{G}/\mathcal{K}),Cl_{\tau}(\mathcal{G}/\mathcal{K})),$
$1_{C_{0}(\mathcal{G}/\mathcal{K})}\otimes_{\mathbb{C}}[\eta_{\mathcal{G}/\mathcal{K}}]\in[\Theta_{\mathcal{G}/\mathcal{K}}].$
Such element $[\eta_{\mathcal{G}/\mathcal{K}}]$ is always represented by a pair $(Cl_{\tau}(\mathcal{G}/\mathcal{K})),c(\eta))$, where $\eta$ is a covector field on $\mathcal{G}/\mathcal{K}$ and $c$ is the Clifford multiplication.
In the case that $\mathcal{G}=Gr(H)_{x}$ for $x\in M$, we see that it is nilpotent and simply-connected, so $\mathcal{K}$ will always be equal to 00.
Definition 3.3.4.
  • There is a continuous (in $x$) family of covector fields $\eta_{Gr(H)_{x}}$ given by the above inductive construction. We denote this family of covector fields by $[\eta_{Gr(H)}]\in\mathcal{R}KK^{Gr(H)}(M;C_{0}(M),Cl_{\tau}(Gr(H)))$. Apply the family of homomorphisms $j^{Gr(H)_{x}}$ to the element $[\eta_{Gr(H)}]$ pointwise for each $x$. The second variable of the resulting $\mathcal{R}KK$-group corresponding to the family of $C^{\ast}$-algebras $\{C^{\ast}(Gr(H)_{x},$ $Cl_{\tau}(Gr(H)_{x}))\}_{x\in M}$, which is isomorphic to $\mathcal{K}(L^{2}(Gr(H)_{x})\otimes Cl_{\tau_{x}})$. By the Morita equivalence between $\mathcal{K}(L^{2}(Gr(H)_{x})$ and $\mathbb{C}$, we obtain an element of the group $\mathcal{R}KK(M;C^{\ast}(Gr(H)),Cl_{\tau}(M))$. We denote it by $[\eta_{C^{\ast}(Gr(H))}]$.
  • For any $x\in M$, consider the Hilbert module $H_{x}=L^{2}(T_{x}^{\ast}M)\otimes Cl_{\tau_{x}}$. The Dirac operator on this Hilbert module is defined by $D_{x}=\sum_{x=1}^{dimM}(-i)c(e_{k})\partial/\partial\xi_{k}$ in any orthonormal basis $\{e_{k}\}$ of $T^{\ast}_{x}M$, where $c(e_{k})$ are the left Clifford multiplication operators and $\xi_{k}$ are the coordinates of $T^{\ast}_{x}M$ in the basis $\{e_{k}\}$. Denote $F_{x}=D_{x}(1+D_{x}^{2})^{-1/2}$ on $H_{x}$. The family of Hilbert modules $H_{x}$, parametrized by $x\in M$, defines a Hilbert module of continuous sections over the algebra $Cl_{\tau}(M)$, vanishing at infinity of $M$, which will be denoted by $\mathcal{E}$. The grading of $\mathcal{E}$ is defined by the grading of the algebra $Cl_{\tau}(M)$. The family of operators $F_{x}$ gives an operator $\Phi$ of degree $1$ on $\mathcal{E}$. The algebra $C_{0}(TM)$ acts on $\mathcal{E}$ on the left by multiplication. We define the fiberwise Dirac element $[d_{\xi}]\in KK^{G}(C_{0}(TM),Cl_{\tau}(M))$ by the pairing $(\mathcal{E},\Phi)$.
  • The Dolbeault element $[D^{\prime}_{M}]\in K^{0}(C^{\ast}(Gr(H)))$ is defined to be the $KK$-product of $[\eta_{C^{\ast}(Gr(H))}]\in\mathcal{R}KK(M;C^{\ast}(Gr(H)),Cl_{\tau}(M))$ and $[d_{M}]\in K^{0}(Cl_{\tau}(M)))$.
Remark 3.3.5.
By Theorem 2.10 in [kasparov2017elliptic], the old Dobeault element $[D_{M}]=[d_{\xi}]\otimes_{Cl_{\tau}(M)}[d_{M}]$, so we see that the new defined Dolbeault element $[D^{\prime}_{M}]$ is taken by replacing $[d_{\xi}]$ with $[\eta_{C^{\ast}(Gr(H))}]$. We can understand that $[\eta_{C^{\ast}(Gr(H))}]$ is the Fourier transform of $[d_{\xi}]$ in some sense.
Now we state the index theorem for H-elliptic operators by Kasparov in [kasparov2024coarsepseudodifferentialcalculusindex].
Definition 3.3.6.
The Clifford cosymbol of an H-elliptic operator $P$ is defined as
$[\sigma_{H}^{cl}(P)]:=[\sigma_{H}(P)]\otimes_{C^{\ast}(Gr(H))}[\eta_{C^{\ast}(Gr(H))}]\in KK(C_{0}(M),Cl_{\tau}(M)).$
Theorem 3.3.7 (Inverse Clifford Index Theorem).
let $P$ be a 00-order properly supported H-elliptic operator on $M$, then
$[\sigma_{H}^{cl}(P)]=[\Theta_{M}]\otimes_{C_{0}(M)}[P]\in KK(C_{0}(M),Cl_{\tau}(M)).$
Theorem 3.3.8 (H-elliptic Index Theorem).
Let $P$ be a 00-order properly supported H-elliptic operator on $M$, then the formula for the index $[P]$ of this operator is
$[P]=[\sigma_{H}(P)]\otimes_{C^{\ast}(Gr(H))}[D^{\prime}_{M}]\in K^{0}(C_{0}(M)).$
Remark 3.3.9.
Comparing with Theorem 3.2.6, we see that this newly defined Dolbeault element $[D^{\prime}_{M}]$ is actually the product of the Connes-Thom element $[\mathcal{E}]$ and the old Dolbeault $[D_{M}]$.